OpenStudy (anonymous):

*Advance Algebra With Financial App ** [Work Is Shown] Bridgette opened a store credit card to purchase a business suit for $463. She put the entire purchase on the credit card. Her APR is 14.99% and the minimum payment is 5%. Assuming she makes only minimum payments, what is the remaining balance on her card after 3 months? -------- 463+14.99%+5%=463.1999 463.1999+14.99%+5%=463.3998 What am I doing wrong ? Can someone help me .

OpenStudy (anonymous):

$396.97 $412.83 $397.35 $445.63

OpenStudy (goldeverything):

$397.35

OpenStudy (anonymous):

How you get that ?

OpenStudy (goldeverything):

just talk the equation out it will make sense

OpenStudy (goldeverything):

and math comes real easy to me so it's hard to explain

jimthompson5910 (jim_thompson5910):

This is something you'll have to step through and simulate Month 1 Monthly Payment = 5% of 463 = 0.05*463 = 23.15 Monthly Payment = $23.15 463*(1+0.1499/12) = 468.783641666667 = 468.78 Old Balance (Before adding on interest): 463 New Balance (After adding on interest): 468.78 Make Monthly Payment: 468.78 - 23.15 = 445.63 ---------------------------------------------------------------------------------------- Month 2 Monthly Payment = 5% of 445.63 = 0.05*445.63 = 22.28 Monthly Payment = $22.28 445.63*(1+0.1499/12) = 451.196661416667 = 451.2 Old Balance (Before adding on interest): 445.63 New Balance (After adding on interest): 451.2 Make Monthly Payment: 451.2 - 22.28 = 428.92 ---------------------------------------------------------------------------------------- Month 3 Monthly Payment = 5% of 428.92 = 0.05*428.92 = 21.45 Monthly Payment = $21.45 428.92*(1+0.1499/12) = 434.277925666667 = 434.28 Old Balance (Before adding on interest): 428.92 New Balance (After adding on interest): 434.28 Make Monthly Payment: 434.28 - 21.45 = 412.83 Remaining balance after month 3 is $412.83